Spiko360:

Solve the equation. (List your answers counterclockwise about the origin starting at the positive real axis.) z8 − i = 0

supie:

\(\bf\ I\ suppose\ we\ are\ solving\ for\ i\) \(z8−i=0\) \(\sf add -z^8\ to\ both\ sides\)\[−i=−z^8\] ÷1 to both sides \(\frac{-i}{-1}=\frac{-z^8}{-1}\) \

supie:

And I don't think we can solve for z... so that's it

lilianamendez2:

yh... u will solve -I/-1=-2^8/-1 to get what I=

lilianamendez2:

and z has no solution

snowflake0531:

@supie 1. i is the imaginary number 2. we can solve for z... we're talking about polar coordinates and forms and that stuff o-0

supie:

oh ok thanx